OpenStudy (anonymous):

what is the square root of 200x^3

OpenStudy (akashdeepdeb):

WELCOME TO OPENSTUDY!! :D

OpenStudy (akashdeepdeb):

Is the expression \[\sqrt{200x^3} ?\] If YES then write it like this. \[\sqrt{200x^3} = \sqrt{10 * 10 * 2 * x * x * x} = 10x \sqrt{2x}\] Understood? :)